OpenStudy (anonymous):

twice a number is 12 less than 6 times the opposite of the number. find the number.

OpenStudy (anonymous):

let the number is x then 2x -12 = -6x

OpenStudy (anonymous):

so do I subtract 2 from 6 then?

OpenStudy (anonymous):

no add 2x and 6x

OpenStudy (anonymous):

that would be -4x since the 6 is negative

OpenStudy (anonymous):

ok let me make it easy for you 2x - 12 = -6x 2x +6x = 12

OpenStudy (anonymous):

now find x

OpenStudy (anonymous):

would that be 8x=12? or am I doing this all wrong..

OpenStudy (johnweldon1993):

And you would see that you set it up incorrectly....all the other steps are correct....however it should have been set up as \[2x = -6x - 12\] From what you are doing...you will get 3/2 correct? when you plug that in to check...it is incorrect....

OpenStudy (anonymous):

okay. thank you

OpenStudy (johnweldon1993):

So....start again from this equation \[2x = -6x - 12\] add 6x to both sides...this will move that -6x to the left side...and make it positive ...so \[8x = -12\] now you divide both sides by 8 \[x = -\frac{ 12 }{ 8 }\] when simplified... \[-\frac{ 3 }{ 2 }\] Now...we check it back in the original equation \[2(-\frac{ 3 }{ 2 }) = -6(-\frac{ 3 }{ 2 }) - 12\] \[-3 = 9 - 12 \] \[-3 = -3\] so -(3/2) would be your answer